# Achieving Human Level Competitive Robot Table Tennis

### v1

@@ I INTRODUCTION
The components described above ultimately lead to competitive gameplay at human level that humans actually enjoy playing with — the resulting policy is flexible, adaptable, extensible, and more interpretable than a monolithic system. In summary, in this paper we introduce the first robot learning system to achieve amateur human level performance in a competitive sport against unseen human opponents demonstrated through a user study. To achieve this goal, we make the following technical contributions (1) a hierarchical and modular policy architecture, specifically the (i) low level controllers with their skill descriptors and (ii) a high level controller that chooses the low level skills; (2) techniques to enable zero-shot sim-to-real including an iterative approach to defining the training task distribution that is grounded in the real-world and defines an automatic curriculum, (3) real time adaptation to unseen opponents and (4) A user-study to test our model playing actual matches against unseen humans in physical environments.
@@ IV-A Robot sports
Since the early days of Artificial Intelligence, Chess and other competitive games have played a critical role in the development of new algorithms and technologies. Examples of highly-skilled AI agents can be found for backgammon [32], chess [33, 34], Go [35], Dota 2 [36], poker [37], and a large number of other games. In many of the these domains, agents have already achieved expert or super human-level performance by playing in a simulated or virtual environment against themselves or a human opponent. A more recent development are competitive games played between people and robotic agents in the real, physical world. In particular robot sports games have gained substantial attention within the research community since they require fast-paced action generation in challenging, non-deterministic and partially-observable environments. A prominent example is the RoboCup competition [38] – a grand robotics challenge with the goal of fielding a fully autonomous robot team that can win a soccer match against the human World Cup champion team by 2050. RoboCup provides a comprehensive research domains that address real-world complexities like sensor fusion, reactive behavior learning, strategic acquisition, planning, multi-agent systems, context recognition, vision, strategic decision-making, motor control, and intelligent robot control in both simulated and hardware environments: e.g.  [39], [40], [41], [42], [43]. Table tennis presents a research field of comparable intricacy, prompting us to devise strategies enabling direct competition with human opponents. Recently [25] demonstrated that RL-trained policies could learn a number of core skills such as tackling, kick and chase, getting up from the ground, and blocking opponent shots in 1 vs 1 robot vs robot humanoid soccer. Other examples for robot sports include fencing [44, 45], skiing [46], tennis [47], catching [48] and badminton [49]. The majority of works and competitions in robot sports focus on either sub-aspects of the game or on playing in substantially simplified settings, rather than realistic and competitive matches with humans.
@@
By contrast, in this paper, we focus playing a competitive game of table tennis against human players of different skill-level and under realistic conditions. Early work on robot table tennis can be traced back to 1983 with Billingsley’s challenge – a competition based on a set of simplified rules that account for technological limitations of the time [4]. Since then, a variety of increasingly skilled table tennis agents have been proposed [50, 51, 52, 53, 9]. A number of works have focused on aspects of action and motion generation, such as returning the ball [5, 54, 55], aiming to a specific target [6, 56], rallying [8, 57, 58], or even smashing [7, 53]. Conversely, other works have focused on aspects of state estimation inherent to the table tennis game, including, for instance, trajectory prediction [59], racket pose detection [60], spin detection [61, 62] or the identification of human strategies [63]. The architecture and framework proposed in this paper leverages a combination of these skills in order to tackle the full competitive game for the first time.
@@ IV-B Sim-to-real
Minimizing the sim-to-real gap for physical dynamics is widely recognized to be a challenging problem and an extensive body of research has been dedicated to closing it in different domains. Multiple techniques have been developed, including system identification [66], learning hybrid simulators [67], dynamics randomization [68] and learning from privileged information [69]. The agent presented in this paper utilises many of these techniques; system ID, dynamics randomization, simulated latency, automatic resets and a policy architecture designed to encourage smooth control. For more details see [9]. Likewise, advances in physics engine development, MuJoCo’s [19] solid state fluid dynamics model for example, also contribute substantially to closing the sim to real gap. In addition, our agent utilizes task-conditioned physics parameters, in which some physical parameters are changed depending on the state of the environment, as well as an additional stage of simulated training to ”spin correct” and our use of sim-to-real adapter layers.
@@ IV-D Hierarchical robot policies
Our work is closest in spirit to the work in [52] in which a gating network is learned to create mixtures of existing low-level policies. The gating network generates probabilities indicating the likelihood that a policy is the right one given the current context. The approach builds on an imitation learning framework and requires expert human demonstrations for each of the lower-level behaviors. In a similar vein, SayCan [90] uses learned value functions in order to determine if a low-level policy will be successful if executed from a given state. However, training SayCan is performed using both supervised and reinforcement learning. The overall objective of SayCan is to use a large language model for high-level planning which is the grounded via value functions and low-level policies in the real world. Our approach deviates from previous methodologies in that it builds upon instance based learning. Skill descriptors of the low-level policies are represented as KD-trees. Hence, new information can incrementally be added without retraining the HLC. We leverage this ability to enable rapid sim to real transfer, i.e., we collect real-world data to improve the accuracy of skill descriptors. In addition, the HLC also continuously collects information about the current opponent, without the need for retraining. As a result, our approach enables real-time learning and is highly adaptive to the current environment and opponent.
@@ V Limitations
Our research demonstrates progress in training robots to play table tennis against human opponents. However, several limitations warrant further investigation: (1) Fast Balls: The robot struggles to react effectively to very fast balls due to multiple factors; lack of data, high system latency, and between-shot reset. Even with a high control frequency (50hz), the inherent latency in the system (${\sim}100ms$) significantly restricts the number of actionable decisions the robot can make within a given timeframe. This problem was exacerbated by the fact that the robot was reset to the same initial pose — close to the table, paddle facing forward — in between each ball hit to account for the single ball training. While increasing the control frequency could mitigate this issue, it reduces the available time per action for computation. (2) High Balls: The robot is unable to handle balls higher than ${\sim}6ft$ above the table, as it goes outside the camera FOV. In practice, it was seldom a problem. (3) Low Balls: The robot’s inability to consistently handle very low balls is primarily attributed to collision avoidance protocols built into the system. These protocols, while essential for protecting the robot’s paddle, pose a significant challenge in handling balls that stay close to the table. This limitation was frequently exploited by skilled players by sending underspin balls that stay very low. (4) Short Pips Paddle: In this work, we chose a paddle that can impart spin, while still being possible to model within the simulator. The paddle doesn’t have the smooth rubber found on advanced table tennis paddles [17]. We found accurately modeling such advanced paddles extremely challenging given the highly nonlinear and multimodal nature of the paddle. This limitation was frequently noted by human participants and highlights the need for more advanced paddle modeling techniques. (5) Spin Detection The policy is unable to accurately read extreme spins on an incoming ball. This limited the amount of spin that could be effectively handled, a fact that was exploited by a number of opponents. It was possible to train specialist LLC policies to handle different types of spin, so improvements in spin detection would likely directly translate to improved overall performance. (6) No Multi-ball Strategic Game Play: The robot’s gameplay is predominantly ”one ball at a time”, lacking the multi-ball strategic play observed in skilled human players. This limitation is attributed to the single-ball training approach. (7) Motion Capture Unreliability: The reliance on motion capture technology introduces a degree of unreliability, as occasional glitches or inaccuracies in motion tracking can negatively impact the robot’s performance. (8) Backhand The policy had weaker backhand play compared with the forehand. (9) Predictability The policy did vary its placement however this could be improved further. Being less predictable in placement of certain balls was one of the top three areas for improvement as identified by the match referee (in addition to reading spin better and improving the backhand).

### v2

@@ I INTRODUCTION
The components described above ultimately lead to competitive gameplay at human level that humans actually enjoy playing with — the resulting policy is flexible, adaptable, extensible, and more interpretable than a monolithic system. In summary, in this paper we introduce the first robot learning system to achieve amateur human level performance in an interactive competitive sport against unseen human opponents demonstrated through a user study. To achieve this goal, we make the following technical contributions (1) a hierarchical and modular policy architecture, specifically the (i) low level controllers with their skill descriptors and (ii) a high level controller that chooses the low level skills; (2) techniques to enable zero-shot sim-to-real including an iterative approach to defining the training task distribution that is grounded in the real-world and defines an automatic curriculum, (3) real time adaptation to unseen opponents and (4) A user-study to test our model playing actual matches against unseen humans in physical environments.
@@ IV-A Robot sports
Since the early days of Artificial Intelligence, Chess and other competitive games have played a critical role in the development of new algorithms and technologies. Examples of highly-skilled AI agents can be found for backgammon [32], chess [33, 34], Go [35], Dota 2 [36], poker [37], and a large number of other games. In many of the these domains, agents have already achieved expert or super human-level performance by playing in a simulated or virtual environment against themselves or a human opponent. A more recent development are competitive games played between people and robotic agents in the real, physical world. In particular, robot sports games have gained substantial attention within the research community since they require fast-paced action generation in challenging, non-deterministic and partially-observable environments. A prominent example is the RoboCup competition [38] – a grand robotics challenge with the goal of fielding a fully autonomous robot team that can win a soccer match against the human World Cup champion team by 2050. RoboCup provides a comprehensive research domain that address real-world complexities like sensor fusion, reactive behavior learning, strategic acquisition, planning, multi-agent systems, context recognition, vision, strategic decision-making, motor control, and intelligent robot control in both simulated and hardware environments: e.g.  [39], [40], [41], [42], [43]. Table tennis presents a research field of comparable intricacy, prompting us to devise strategies enabling direct competition with human opponents. Recently [25] demonstrated that RL-trained policies could learn a number of core soccer skills such as tackling, kick and chase, getting up from the ground, and blocking opponent shots in 1 vs. 1 robot vs. robot humanoid soccer. Other examples for robot sports include fencing [44, 45], skiing [46], tennis [47], catching [48] and badminton [49]. The majority of works and competitions in robot sports focus on either sub-aspects of the game or on playing in substantially simplified settings, rather than realistic and competitive matches with humans. There is one notable exception [50], which demonstrated impressive champion level performance in drone racing on a fixed track known in advance in a head-to-head race vs. human experts. By contrast, this work focuses on an interactive sport in which the robot must continuously respond to an unseen human’s actions instead of a head-to-head time trial.
@@
In this paper, we focus playing a competitive game of table tennis against human players of different skill-levels and under realistic conditions. Early work on robot table tennis can be traced back to 1983 with Billingsley’s challenge – a competition based on a set of simplified rules that account for technological limitations of the time [4]. Since then, a variety of increasingly skilled table tennis agents have been proposed [51, 52, 53, 54, 55, 56, 57, 9]. A number of works have focused on aspects of action and motion generation, such as returning the ball [58, 5, 59, 60, 61], aiming to a specific target [6, 62], rallying [8, 63, 64], or even smashing [7, 57]. Conversely, other works have focused on aspects of state estimation inherent to the table tennis game, including, for instance, trajectory prediction [65], racket pose detection [66], spin detection [67, 68] or the identification of human strategies [69, 70, 71]. The architecture and framework proposed in this paper leverages a combination of these skills in order to tackle the full competitive game for the first time.
@@ IV-B Sim-to-real
Minimizing the sim-to-real gap for physical dynamics is widely recognized to be a challenging problem and an extensive body of research has been dedicated to closing it in different domains. Multiple techniques have been developed, including system identification [74], learning hybrid simulators [75], dynamics randomization [76] and learning from privileged information [77]. The agent presented in this paper utilises many of these techniques; system ID, dynamics randomization, simulated latency, automatic resets and a policy architecture designed to encourage smooth control. For more details see [9]. Likewise, advances in physics engine development, MuJoCo’s [19] solid state fluid dynamics model for example, also contribute substantially to closing the sim to real gap. In addition, our agent utilizes task-conditioned physics parameters, in which some physical parameters are changed depending on the state of the environment, as well as an additional stage of simulated training to “spin correct” and our use of sim-to-real adapter layers.
@@ IV-D Hierarchical robot policies
A much earlier version of our system [98] demonstrated the effectiveness of a hierarchical approach to segment strategy, targeting, and execution for simulated robot table tennis, with a major difference being that the lower levels were model-based and were thus difficult to translate from simulation to the real world.
Our current work is closest in spirit to the work in [55] in which a gating network is learned to create mixtures of existing low-level policies. The gating network generates probabilities indicating the likelihood that a policy is the right one given the current context. The approach builds on an imitation learning framework and requires expert human demonstrations for each of the lower-level behaviors. In a similar vein, SayCan [99] uses learned value functions in order to determine if a low-level policy will be successful if executed from a given state. However, training SayCan is performed using both supervised and reinforcement learning. The overall objective of SayCan is to use a large language model for high-level planning which is the grounded via value functions and low-level policies in the real world. Our approach deviates from previous methodologies in that it builds upon instance based learning. Skill descriptors of the low-level policies are represented as KD-trees. Hence, new information can incrementally be added without retraining the HLC. We leverage this ability to enable rapid sim to real transfer, i.e., we collect real-world data to improve the accuracy of skill descriptors. In addition, the HLC also continuously collects information about the current opponent, without the need for retraining. As a result, our approach enables real-time learning and is highly adaptive to the current environment and opponent.
@@ V Limitations
Our research demonstrates progress in training robots to play table tennis against human opponents. However, several limitations warrant further investigation: (1) Fast Balls: The robot struggles to react effectively to very fast balls due to multiple factors; lack of data, high system latency, and between-shot reset. Even with a high control frequency (50hz), the inherent latency in the system (${\sim}100ms$) significantly restricts the number of actionable decisions the robot can make within a given timeframe. This problem was exacerbated by the fact that the robot was reset to the same initial pose — close to the table, paddle facing forward — in between each ball hit to account for the single ball training. While increasing the control frequency could mitigate this issue, it reduces the available time per action for computation. (2) High Balls: The robot is unable to handle balls higher than ${\sim}6ft$ above the table, as it goes outside the camera FOV. In practice, it was seldom a problem. (3) Low Balls: The robot’s inability to consistently handle very low balls is primarily attributed to collision avoidance protocols built into the system. These protocols, while essential for protecting the robot’s paddle, pose a significant challenge in handling balls that stay close to the table. This limitation was frequently exploited by skilled players by sending underspin balls that stay very low. (4) Short Pips Paddle: In this work, we chose a paddle that can impart spin, while still being possible to model within the simulator. The paddle doesn’t have the smooth rubber found on advanced table tennis paddles [17]. We found accurately modeling such advanced paddles extremely challenging given the highly nonlinear and multimodal nature of the paddle. This limitation was frequently noted by human participants and highlights the need for more advanced paddle modeling techniques. (5) Spin Detection: The policy is unable to accurately read extreme spins on an incoming ball. This limited the amount of spin that could be effectively handled, a fact that was exploited by a number of opponents. It was possible to train specialist LLC policies to handle different types of spin, so improvements in spin detection would likely directly translate to improved overall performance. (6) No Multi-ball Strategic Game Play: The robot’s gameplay is predominantly ”one ball at a time”, lacking the multi-ball strategic play observed in skilled human players. This limitation is attributed to the single-ball training approach. (7) Motion Capture Unreliability: The reliance on motion capture technology introduces a degree of unreliability, as occasional glitches or inaccuracies in motion tracking can negatively impact the robot’s performance. (8) Backhand The policy had weaker backhand play compared with the forehand. (9) Predictability The policy did vary its placement however this could be improved further. Being less predictable in placement of certain balls was one of the top three areas for improvement as identified by the match referee (in addition to reading spin better and improving the backhand).